### v2

@@ 4.1 Currents and Fields
In the limit of large r, the enclosed current approaches the total current:
$I_{\theta}(\infty)=-\frac{\pi^{3/2}B_{0}\sigma}{\mu_{0}}$ (19)
Figure 7 shows an overlay of the total current, current density, and cyclotron magnetic field as a function of the cylinder radius. It is seen there that the total local current encompassing the cylinder is on the order of $1.3\times 10^{6}$ amperes.
@@ 5 Rigid Rotor Model
We begin with Ohm’s law for the azimuthal electric field as discussed by Breginskii and Bettencourt for a magnetised plasma Braginskii (1965); Bittencourt (2013):
@@
and substituting $E_{\theta}(r)$ and $B_{z}(r)$ into the expression for $\omega_{r}(r)$:
$\omega_{r}(r)$ $=\frac{-\frac{2B_{0}}{\mu_{0}\sigma_{e}\sigma^{2}}r\exp\left(-\frac{r^{2}}{\sigma^{2}}\right)}{rB_{0}\exp\left(-\frac{r^{2}}{\sigma^{2}}\right)}$ (25)
$=-\frac{2}{\mu_{0}\sigma_{e}\sigma^{2}}$ (26)
It is seen that $r$, $B_{0}$, and exponential terms cancel, leaving a constant:
@@ 6 Resonant Isotope Enrichment- Ion Bernstein Waves
The anomalously high isotope entrenchments for specific charge states, as seen in Figure 3 (and to a lesser extent in Figure 2), are suggestive of some kind of resonance enhancement of the enrichment process. This effect can be observed for example in Fig.3 for the specific charge states of Ni-58 at charge states $Z=9,8,6$ , and in particular being more dominant at smaller expansion angles. This behavior is consistent with and is, we propose, caused by resonant coupling between the cyclotron rotating ions and associated ion Bernstein waves (IBWs) that exist within a magnetized plasma Boyd and Sanderson (2003); Bittencourt (2013). IBWs are longitudinal electrostatic waves that propagate perpendicular to the magnetic field and exist in harmonics of the ion cyclotron frequency. These waves can interact resonantly with ions when their wave frequency matches the ion’s cyclotron frequency or one of its harmonics. In our ultrafast laser ablation plasma, the strong, self-generated magnetic field produced in the early expansion phase enables a localized IBW spectrum that overlaps with the higher cyclotron frequencies of high charge state ions. Because these ions are also among the fastest particles in the plasma, they traverse the high-field region rapidly and are preferentially positioned to undergo coherent wave-particle coupling. This coupling leads to selective energy transfer and preferential transport of certain isotopes—such as Ni-58—whose mass-to-charge ratio allows for near-exact resonance with the available IBW modes. The sharpness of these enhancements, and their confinement to specific charge states and angular observation regions, cannot be easily explained by other mechanisms such as thermal diffusion, bulk drift rotation, or other hydrodynamic processes . Instead, they point to a nonlinear, field-mediated resonance mechanism such as IBW coupling, which can selectively amplify isotope populations far beyond what would be expected from normal equilibrium behavior. The presence of IBW-driven enrichment signatures, especially in low-density, high-$Z$ ion populations, thus serves as a compelling indicator of localized, resonance-based separation dynamics acting within the early-time magnetic structure of the plasma.
@@ 6.2 Increasing the Gaussian Width to Match Observed Resonance Positions
In our model, as presented above, the magnetic field profile is described by a Gaussian where $\sigma_{r}$ characterizes the spatial extent of the magnetic field within the plasma. The resonance condition for each charge state is set by matching the ion’s cyclotron frequency to a localized IBW, yielding the resonance radius of Eq. 32. This relation states that the resonance radius scales directly with $\sigma_{r}$. However, when applying this model to experimental data using the initially determined field width (from Eq. 8, $\sigma_{r}=56\,\mu\text{m}$), the predicted resonance positions for lower charge states (e.g., $Z=8,7,6$) fall systematically short of the observed angular locations by a factor of about 3. To bring the calculated resonance positions into agreement with experimental enrichment angles, the Gaussian width must be increased from $\sigma_{r}=56\,\mu\text{m}$ to $\sigma_{r}=182\,\mu\text{m}$ . Physically, this corresponds to a broader magnetic field distribution that drops off more slowly with radius, allowing lower charge states to resonate farther from the center which is where they are actually observed. Increasing $\sigma_{r}$ in this way does not require a higher peak field $B_{0}$, but instead simply reflects a magnetic field that more uniformly fills the centrifuge cylinder providing a broader spatial variation of the magnetic field with radius. This field density relaxes, through dissipative effects in the plasma, to evolve into the more narrow and concentrated quasi-equilibrium field that effects the observed bulk isotope enrichment. See Fig.8 for a comparison of the shapes of these two fields.
@@
These ions are more abundant and have lower energies ($\approx 0.2$ keV), resulting in slower transit times of a few nanoseconds. They experience the relaxed magnetic field structure, which governs the bulk isotope separation. This sets the upper and lower bounds for the magnetic field creation and collapse time scales. Using the energy vs charge state data of Fig.10 (in Appendix section 8.1) for velocity information, the early magnetic field with wide gaussian must form on a very short time scale, significantly shorter than 537 picoseconds, which is the 125 micron transit time of the fastest +9 ion at 16.3 keV. This early field needs to last for a time of 835 picoseconds which is the transit time of the charge state +6 ion at 6 keV which is observed resonating at 60 degrees and is the slowest of the resonating ions. The final quasi-stationary field of the 56 micron Gaussian width needs to last for a time that is at least 4.85 nanoseconds to clear the transit time of the slowest low energy +1 ions at approximately 0.2 keV..
@@
While high charge state ions (e.g., $Z=9,8,6$) exhibit strong and selective resonant enhancements in our experimental data, it is important to recognize that these ions represent a very small fraction of the total plasma population being three to four orders of magnitude lower in density than the dominant low charge states ($Z=1$ and $Z=2$). Consequently, although the high-$Z$ species provide clear evidence of coherent wave-particle interactions, such as coupling to ion Bernstein waves (IBWs), they are unlikely to drive the bulk of the observed isotope separation. Instead, the primary separation process is governed by the dynamics of the much more abundant low charge state ions, which respond to the overall plasma structure. In this context, the high-$Z$ resonant enhancements serve primarily as diagnostic tracers of underlying plasma wave activity and magnetic field structure during the early stages of plasma evolution. They provide key information about the plasma’s resonant behavior without being directly responsible for the majority of mass separation observed at later quasi-equilibrium stages. It is instructive to think about how these resonances could be brought down to the low Z ions in order to enhance energy coupling to them and the bulk of the mass enrichment process. It would require much higher magnetic fields to accomplish this along with possibly using very low mass elements such as lithium, beryllium, or boron since cyclotron frequencies scale linearly with B and with mass as 1/m.

### v4

@@ 4.1 Currents and Fields
In the limit of large r, the enclosed current approaches the total effective current where $B_{0}$ is $B_{eff}$ at r = 0 and
$I_{eff}=-\frac{\pi^{3/2}B_{0}\sigma}{\mu_{0}}$ (19)
Figure 7 shows an overlay of the total current, current density, and cyclotron effective magnetic field as a function of the cylinder radius. It is seen there that the total effective current encompassing the cylinder is on the order of $1.3\times 10^{6}$ amperes. This value will be further broken down in Section 7 where $B_{eff}$ is discussed.
@@ 5 Rigid Rotor Model
We begin with Ohm’s law for the azimuthal electric field as discussed by Breginskii and Bettencourt for a magnetised plasma braginskii1965transport; bittencourt2013fundamentals:
@@
and substituting $E_{\theta}(r)$ and $B_{z}(r)$ into the expression for $\omega_{r}(r)$ it is seen that $r$, $B_{0}$, and exponential terms cancel, leaving a constant:
@@ 6 Resonant Isotope Enrichment- Ion Bernstein Waves
The anomalously high isotope entrenchments for specific charge states, as seen in Figure 3 (and to a lesser extent in Figure 2), are suggestive of some kind of resonance enhancement of the enrichment process. This effect can be observed for example in Fig.3 for the specific charge states of Ni-58 at charge states $Z=9,8,6$ , and in particular being more dominant at smaller expansion angles. This behavior is consistent with and is, we propose, caused by resonant coupling between the cyclotron rotating ions and associated ion Bernstein waves (IBWs) that exist within a magnetized plasma boyd2003physics; bittencourt2013fundamentals. IBWs are longitudinal electrostatic waves that propagate perpendicular to the magnetic field and exist in harmonics of the ion cyclotron frequency. These waves can interact resonantly with ions when their wave frequency matches the ion’s cyclotron frequency or one of its harmonics. In our ultrafast laser ablation plasma, the strong, self-generated magnetic field produced in the early expansion phase enables a localized IBW spectrum that overlaps with the higher cyclotron frequencies of high charge state ions. Because these ions are also among the fastest particles in the plasma, they traverse the high-field region rapidly and are preferentially positioned to undergo coherent wave-particle coupling. This coupling leads to selective energy transfer and preferential transport of certain isotopes—such as Ni-58—whose mass-to-charge ratio allows for near-exact resonance with the available IBW modes. The sharpness of these enhancements, and their confinement to specific charge states and angular observation regions, cannot be easily explained by other mechanisms such as thermal diffusion, bulk drift rotation, or other hydrodynamic processes . Instead, they point to a nonlinear, field-mediated resonance mechanism such as IBW coupling, which can selectively amplify isotope populations far beyond what would be expected from normal equilibrium behavior. The presence of IBW-driven enrichment signatures, especially in low-density, high-$Z$ ion populations, thus serves as a compelling indicator of localized, resonance-based separation dynamics acting within the early-time magnetic structure of the plasma.
@@ 6.2 Increasing the Gaussian Width to Match Observed Resonance Positions
In our model, as presented above, the magnetic field profile is described by a Gaussian where $\sigma_{r}$ characterizes the spatial extent of the magnetic field within the plasma. The resonance condition for each charge state is set by matching the ion’s cyclotron frequency to a localized IBW, yielding the resonance radius of Eq. 32. This relation states that the resonance radius scales directly with $\sigma_{r}$. However, when applying this model to experimental data using the initially determined field width (from Eq. 8, $\sigma_{r}=56\,\mu\text{m}$), the predicted resonance positions for lower charge states (e.g., $Z=8,7,6$) fall systematically short of the observed angular locations by a factor of about 3. To bring the calculated resonance positions into agreement with experimental enrichment angles, the Gaussian width must be increased from $\sigma_{r}=56\,\mu\text{m}$ to $\sigma_{r}=182\,\mu\text{m}$ . Physically, this corresponds to a broader magnetic field distribution that drops off more slowly with radius, allowing lower charge states to resonate farther from the center which is where they are actually observed.
@@
Justification and support for collapsing the early field distribution to a narrower one can be found in the work of bell1993observation where they demonstrate, with 2.1 picosecond near-infrared pulses, an azimuthal magnetic pinch effect collimating the plasma jet and focusing the ions into a half angle of 10 degrees about the normal to the ablating surface after laser pulse plasma formation. They demonstrate this both experimentally with time integrated x-ray images as well as computationally with a magneto hydrodynamic code demonstrating temperature and magnetic field contours about the normal direction with collimated jets of 100-to-240-micron length. Visual demonstration of such jet formation can also be found in albert2003time. In the present case therefore, the core of the plasma is retained intact, with the pinched peak field remaining the same, while the outer edges dissipate. Increasing $\sigma_{r}$ in this way does not require a higher peak field $B_{0}$, but instead simply reflects a magnetic field that more uniformly fills the centrifuge cylinder, prior to full pinching, providing a broader spatial variation of the magnetic field with radius. This initial field thereafter relaxes, through dissipative effects in the outer wings of the plasma, to evolve into the more narrow and concentrated quasi-equilibrium field that is responsible for the observed bulk isotope enrichment. See Fig.8 for a comparison of the shapes of these two fields. Further demonstration of such pinching can be found in borghesi1998megagauss
@@
These slower ions are more abundant and have lower energies ($\approx 0.2$ keV), resulting in longer transit times of a few nanoseconds. They experience the relaxed magnetic field structure, which governs the bulk isotope separation.This sets the upper and lower bounds for the magnetic field creation and collapse time scales. Using the energy vs charge state data of Fig.10 (and Appendix section 8.1) for velocity information, the early magnetic field with wide gaussian must form on a very short time scale, significantly shorter than 537 picoseconds, which is the 125 micron transit time of the fastest +9 ion at 16.3 keV. This early field needs to last for a time of 835 picoseconds which is the transit time of the charge state +6 ion at 6 keV which is observed resonating at 60 degrees and is the slowest of the resonating ions. The final quasi-stationary field of the 56 micron Gaussian width needs to last for a time that is at least 4.85 nanoseconds to clear the transit time of the slowest low energy +1 ions at approximately 0.2 keV.
@@
While high charge state ions (e.g., $Z=9,8,6$) exhibit strong and selective resonant enhancements in our experimental data, it is important to recognize that these ions represent a very small fraction of the total plasma population being three to four orders of magnitude lower in density than the dominant low charge states ($Z=1$ and $Z=2$). Consequently, although the high-$Z$ species provide clear evidence of coherent wave-particle interactions, such as coupling to ion Bernstein waves (IBWs), they are unlikely to drive the bulk of the observed isotope separation. Instead, the primary separation process is governed by the dynamics of the much more abundant low charge state ions, which respond to the overall plasma structure. In this context, the high-$Z$ resonant enhancements serve primarily as diagnostic tracers of underlying plasma wave activity and magnetic field structure during the early stages of plasma evolution. They provide key information about the plasma’s resonant behavior without being directly responsible for the majority of mass separation observed at later quasi-equilibrium stages. It is instructive to think about how these resonances could be brought down to the low Z ions in order to enhance energy coupling to them and the bulk of the mass enrichment process. Changing laser pulse parameters and plasma species suggests itself as possible ways to accomplish that. See Figures 6 and 8 in reference vanrompay2000isotope as an example, where laser pulse durations are on the order of the cyclotron rotation period for Boron isotopes.
7 Broad Spectrum IBW and $B_{\mathrm{eff}}$
Having established the presence of IBW resonance waves in our ablation plasma, it is instructive to extend their role to the more generalized broad spectral case beyond the resonances of the previous section. Focusing on the late-time stable field condition and recognizing that since there is, in the plasma, a range of ion energies with different charge states, along with a radially varying magnetic field, there will exist a broad spectrum of cyclotron orbits with a similarly structured spectrum of IBW waves. As such, these spectral waves can drive near-resonant pumping across the entire distribution of ion cyclotron orbits. They will thereby produce larger values in the separation equation (Eq. 2) beyond that expected from the magnetic field rotation alone. Exact resonance isn’t required; proximity within the finite linewidth of each harmonic is enough. This gives a global, yet charge-selective, perpendicular heating and will greatly amplify isotope separation beyond the static B-field expectation alone.
As described in the previous section, an IBW wave is a longitudinal electrostatic wave whose k vector is perpendicular to the magnetic field and therefore aligned to transfer energy to the cyclotron orbits when sufficient frequency overlap occurs. These pulses of energy will increase the perpendicular orbital velocity by a delta increment upon each damping interaction. These velocity increments lead to an increased radial increment on the cyclotron orbits which results in an increase in isotopic fractionation.
Looking at the centrifuge equation (Eq. 1), one sees that the enrichment ratio is exponentially dependent on both the rotation rate and the radius of rotation. This means there are two different channels by which enrichment can occur. The magnetic field controls the cyclotron ion rotation rate whereas the transverse velocity controls their radius of rotation thereby resulting in two distinct enrichment mechanisms, one contingent on the existence of the other. If we assume only the B field is responsible for separations, when in fact an incremental radial contribution is also at play, then the extracted value of the B field will appear larger than its contribution is actually providing. This would appear to be the case for our originally extracted B field in Section 3.2.
By way of example, consider the centrifuge plasma separation as being essentially a Boltzmann-in-a-rotating-frame, being expressed here as a sample case s, in a generalized cylinder of radius r as:
$n_{s}(r)=n_{s}(0)\,\exp\!\left[\tfrac{\frac{1}{2}\Delta m_{s}\omega_{r}^{2}r^{2}}{T_{s}}\right].$
This relation shows a strong dependence on $r^{2}$ through the centrifugal term $\tfrac{1}{2}\Delta m_{s}\omega_{r}^{2}r^{2}$. Now suppose an (IBW) resonance pushes just one species outward by a small radial increment $\Delta r$. Then the ratio of densities becomes
$\frac{n_{s}(r+\Delta r)}{n_{s}(r)}\approx\exp\!\left[\frac{\Delta m_{s}\omega_{r}^{2}}{2T_{s}}\left(2r\,\Delta r+\Delta r^{2}\right)\right].$ (32)
If the background rotation $\omega_{r}$ is large (as in our centrifuge case), even a modest $\Delta r$ yields a large exponential factor. This explains the extraordinary separation values and large effective B field previously extracted. The IBWs only need to slightly nudge the broad spectrum of near-resonant species outward to achieve strong enrichment.
To this end, the extracted magnetic field B can be expressed as a linear combination of the actual longitudinal $B_{z}$ plus an additional electrostatic $B_{ibw}$ to account for the total extracted B effective given as $B_{eff}=B_{z}+B_{ibw}$.
Although Ion Bernstein Waves (IBWs) are electrostatic modes and do not contribute directly to the quasi-static longitudinal magnetic field $B_{z}$, their presence profoundly alters the effective rotational dynamics inferred from the isotope-enrichment data. In our system, the centrifuge rotation is driven by the true axial field $B_{z}$, but the IBW spectrum—generated by mode conversion of electron plasma waves into ion Bernstein harmonics—produces additional azimuthal velocity and finite-radius excursions in the ions. These oscillatory IBW fields occur near integer multiples of the ion cyclotron frequency and therefore mimic the structure of a cyclotron spectrum, even though they do not supply additional axial magnetic flux. When the resulting ion distributions are interpreted using the standard centrifuge Boltzmann formalism, the combined cyclotron motion and IBW-driven excursions will appear as an enhanced effective exponential rotation term $\Omega_{\mathrm{eff}}^{2}=\Omega_{\mathrm{ci}}^{2}+\Omega_{\mathrm{IBW}}^{2}$ where $\Omega_{\mathrm{ci}}^{2}$ is the ion cyclotron rotation rate and $\Omega_{\mathrm{IBW}}^{2}$ is the IBW contribution to the total effective rotation term. These two rotation contributions are added in quadrature since they appear as energy components in the Boltzmann exponent which requires the use of mean squared rotation factors. In this regard, $\Omega_{\mathrm{eff}}$, can be understood as the result of an “effective” magnetic field
$B_{\mathrm{eff}}\equiv\frac{m}{q}\,\Omega_{\mathrm{eff}}.$
where $B_{\mathrm{eff}}=B_{z}+B_{\mathrm{IBW}}$, not to imply the presence of a second magneto-static field, but as a convenient book-keeping device: $B_{\mathrm{IBW}}$ denotes the equivalent magnetic-field strength that would produce, via ordinary cyclotron motion, the same additional mean-square azimuthal velocity and radial displacement that are actually produced electrostatically by the IBWs. This representation preserves the familiar centrifuge formalism while faithfully incorporating the wave–particle dynamics responsible for the observed enhancements in $\Omega_{\mathrm{eff}}$.
7.1 B Component Contributions
It is constructive to now examine the way in which these two separation components contribute and interact. This is conveniently done by examining how different charge states contribute to the observed rotations within the plasma. For convenience, let us decompose the ”effective” rotation frequency obtained from the original enrichment data using the system average taken over all detector observation angles. That hybrid rotation rate taken from Sec. 3.1 was
$\Omega_{\mathrm{eff}}=3.2\times 10^{9}\ \mathrm{rad/s},$
Breaking it into two parts: (1) the true cyclotron rotation driven by the longitudinal magnetic field $B_{z}$, and (2) an additional contribution associated with the ion Bernstein wave (IBW) spectrum,
$\Omega_{\mathrm{eff}}^{2}=\Omega_{\mathrm{ci}}^{2}+\Omega_{\mathrm{IBW}}^{2}$ (33)
the decomposition is defined by
$\Omega_{\mathrm{ci}}=\frac{ZeB_{z}}{m},\qquad\Omega_{\mathrm{IBW}}=\sqrt{\Omega_{\mathrm{eff}}^{2}-\Omega_{\mathrm{ci}}^{2}},$ (34)
which yields a real IBW term whenever $\Omega_{\mathrm{ci}}\leq\Omega_{\mathrm{eff}}$. This provides a practical criterion for determining which charge states can “participate” in generating the observed effective rotation and at what longitudional field strength.
By way of reference, it is seen from Figure 3 that charge state 10 participates in the IBW resonance at 15 degrees. Thus, we can take that as the highest Z value present in the IBW affected plasma for our decomposition analysis. Using the equalities in Equation 34, the decomposition was performed for charge states $Z=1$ through $Z=10$ using various longitudinal magnetic field values for $B_{z}$. It was found that $B_{z}$ cannot exceed 1.9 MG in order for $\Omega_{\mathrm{eff}}$ to remain real.
Z $\boldsymbol{\Omega_{\mathrm{ci}}}$ (rad/s) $\boldsymbol{\Omega_{\mathrm{IBW}}}$ (rad/s)
1 $3.16\times 10^{8}$ $3.18\times 10^{9}$
2 $6.32\times 10^{8}$ $3.14\times 10^{9}$
3 $9.48\times 10^{8}$ $3.06\times 10^{9}$
4 $1.26\times 10^{9}$ $2.94\times 10^{9}$
5 $1.58\times 10^{9}$ $2.78\times 10^{9}$
6 $1.90\times 10^{9}$ $2.58\times 10^{9}$
7 $2.21\times 10^{9}$ $2.31\times 10^{9}$
8 $2.53\times 10^{9}$ $1.96\times 10^{9}$
9 $2.84\times 10^{9}$ $1.47\times 10^{9}$
10 $3.16\times 10^{9}$ $5.02\times 10^{8}$
Table 4: Cyclotron and IBW rotation frequencies for maximum $B_{z}=1.9~\mathrm{MG}$
The results, shown in Table 4, demonstrate several important trends. First, for this value of $B_{z}$, all charge states produce real IBW contributions, since $\Omega_{\mathrm{cf}}$ remains below $\Omega_{\mathrm{eff}}$ for the entire range $Z=1$ to $Z=10$. This means that every charge state examined contributes through some combination of cyclotron and IBW-driven motion.
Second, because $\Omega_{\mathrm{ci}}\propto Z$, the lower charge states have very small cyclotron frequencies and therefore require substantial IBW contributions to reach the observed rotation rate. For example, at $Z=1$ the cyclotron frequency is only $3.2\times 10^{8}\ \mathrm{rad/s}$, nearly an order of magnitude smaller than $\Omega_{\mathrm{eff}}$, so the IBW term supplies almost all the observed rotation for this species.
Third, as the charge state increases, the cyclotron term becomes a larger fraction of the total rotation. By $Z=10$ the cyclotron frequency has risen to $3.16\times 10^{9}\ \mathrm{rad/s}$, approaching the observed effective frequency. The corresponding IBW frequency is smaller but still nonzero, indicating that even high charge states experience wave-driven motion.
Overall, the pattern follows directly from the relations in Equation 34. As $Z$ increases, the cyclotron term grows linearly while the IBW term decreases accordingly, yet for $B_{z}=1.9~\mathrm{MG}$ the IBW contribution remains real for all charge states considered. This result demonstrates that, under these conditions, all charge states from $Z=1$ to $Z=10$ can participate in generating the observed rotation through a combined action of true cyclotron motion and IBW-induced transverse dynamics. The lower charge states are dominated by the IBW component, while the higher charge states derive more of their rotation from the magnetic field but still retain a meaningful IBW contribution.
By way of completing the discussion on these combined enrichment channels, it should be mentioned that if, instead of using the generic system wide single frequency for demonstrating Eq. 34, we use the observed angularly dependent rotation rates from Table 1 for Ni isotopes, the overall outcome will be the same except that the results would produce a radially changing value of $B_{z}$ reflecting the Gaussian like nature of the data points and, as shown in Table 5 for the Nickel isotopes. The numerical average here is 3.35 MG.
$\theta$ (deg) $B_{z,\max}$ (MG)
$0^{\circ}$ 7.6
$15^{\circ}$ 3.9
$30^{\circ}$ 1.5
$60^{\circ}$ 0.41
Table 5: Maximum longitudinal field consistent with a real IBW contribution at each observation angle for $\Omega_{eff}$ taken from Table 1 for Ni isotopes.
There are two relevant sets of published reports that can be drawn upon for comparison to our experiment and extracted values of longitudinal magnetic field $B_{z}$. briand1985axial; sandhu2002laser; sandhu2001femtosecond. The first involves normal incidence laser irradiation and the second oblique incidence similar to the setup in this report. It can be argued that oblique incidence will produce a more pronounced longitudinal field compared to normal incidence due to tangential components of laser E field at the surface. Brand et al, using Faraday rotation and normal incidence, measure a longitudinal field of 0.6 MG in aluminum with a predicted theoretical maximum of 2.5 MG. Sandhu et al measure ellipticity changes of a reflected probe pulse on aluminum and report a field with peak value of 27 MG droping to 5 MG over a 2 to 10 picoseconds time scale. Their B field is considered perpendicular to the probe propagation k vector. As such, due to oblique geometry, their determined B value will consist of both toroidal and longitudinal components. The relative contribution of each cannot be determined from their experiment, however their B field approaches 5 MG at long times of which 2.5 MG could be longitudinal. Likewise, they also report a similar experiment on copper with asymptotic B fields of 4 MG of which 2 MG could be longitudinal. It can be concluded that our results for the longitudinal $B_{z}$ field are consistent with direct measurements of magnetic field in experimental conditions similar to ours.